OpenStudy (anonymous):

Integrate (please show steps!): e^-5x

OpenStudy (anonymous):

we know that \[\frac{d}{dx} e^{kx} = ke^{kx}\] so doing this in reverse we get \[\int\limits{e^{kx} }dx = \frac{e^{kx}}{k}\]

OpenStudy (anonymous):

+ c of course

OpenStudy (anonymous):

e^-5x dx e^-5x / -5 +c so it is -1/5*e^-5x + C

OpenStudy (kropot72):

First assume that e^5x is the result of integration. Then differentiate e^-5x and obtain -5 * e^5x. So if the trial result is divided by -5 we have the solution which is: (e^-5x)/-5

OpenStudy (kropot72):

Plus C the constant of integration of course.